# Fractal and Regular Geometry of Deep Neural Networks

### v1

@@ 5.1   Proof of item i)
$\lim_{k\to 0}\lim_{h\to 0}\mathbb{E}\left[\frac{f(s+he_{i})-f(s)}{h}\cdot\frac{f(s+ke_{i})-f(s)}{k}\right]\ .$ (5.4)
@@
$\Sigma_{f}(s,t)=\mathbb{E}[T(\varphi^{-1}(s))T(\varphi^{-1}(t))]=\kappa\left(1-\frac{2\|s-t\|^{2}}{(1+\|s\|^{2})(1+\|t\|^{2})}\right)\ .$ (5.5)
Instead of (5.4), we compute a more general limit (which will also be useful for the next steps). Define
$A_{h,k}(s,t;i,j)=\mathbb{E}\left[\frac{f(t+he_{i})-f(t)}{h}\cdot\frac{f(s+ke_{j})-f(s)}{k}\right]\ ,$
and note that (5.4) corresponds to $\lim_{k\to 0}\lim_{h\to 0}A_{h,k}(s,s;i,i)$. Set
@@ 5.2   Proof of item ii)
The second term in (5.9) can be written, using Lemma 5.2 and (5.8), as
@@
Finally to compute the third term in (5.9) one can use Lemma 5.2, eq. 5.7 and the formula for $\kappa^{\prime}$ obtained by deriving (3.1) to get that
@@ 5.3   Proof of item iii)
as a consequence, taking $k=d$ we have $2d\rho(\beta)\leq d+d$ and therefore (5.3) holds with $\zeta_{1}=2d$, $\zeta_{2}=d$ and $K=C^{k}(2k-1)!!$!.
From (5.2), we can compute $\mathbb{E}[D^{i}f(s)^{2}]$ and $\mathbb{E}[D^{i}f(t)^{2}]$, whence to estimate (5.10) we only need to evaluate $\mathbb{E}[D^{i}f(s)D^{i}f(t)]$. Using again Lemma 5.2 we obtain
@@
Since $21<\beta<$ 2 we have also $\|t-s\|^{2}=o(\|t-s\|^{2\beta-2})$ and therefore
@@
6   Proof of Theorem 3.12: boundary volumes in the Kac-Rice class
In Proposition 6.2, we compute the expected values of the boundary volumes of a $C^{1}$ isotropic Gaussian random field. Theorem 3.12 follows using the fact that $\kappa_{L}^{\prime}(1)=\kappa^{\prime}(1)^{L}$ as proved in  [24]. We note that, using Proposition 3.14, under the assumptions of Theorem 3.12, the fields $T_{L}$ are $C^{1}$ a.s for every $L$.
In the proof of Proposition 6.2, we use the following result that can be obtained combining Theorem .2 and Theorem 2.2 in [5].
@@
$\mathbb{E}[\mathcal{H}_{d-1}(A_{u}(X))]=\int_{\mathbb{S}^{d}}\mathbb{E}\left.\big{[}\|\nabla X(t)\|\,\right|\,X(t)=u\big{]}p_{X(t)}(u)\operatorname{dVol}(t)\ ,$
@@
$\mathbb{E}\left[\mathcal{H}_{d-1}(A_{u}(T))\right]=\omega_{d-1}\kappa^{\prime}(1)^{1/2}e^{-u^{2}/2}\qquad a.s.\ .$
@@
Theorem 6.1 implies
$\mathbb{E}\left[\mathcal{H}_{d-1}(A_{u}(T))\right]=\int_{\mathbb{S}^{d}}\mathbb{E}\left[\sqrt{\nabla T(x)^{t}\nabla T(x)}\;|\;T(x)=u\right]p_{T(x)}(u)\operatorname{dVol}(x)\ .$
@@
$\mathbb{E}\left[\mathcal{H}_{d-1}(A_{u}(T))\right]=\frac{1}{\sqrt{2\pi}}e^{-{u^{2}}/2}\int_{\mathbb{S}^{d}}\mathbb{E}\left[\sqrt{\nabla T(x)^{t}\nabla T(x)}\;|\;T(x)=u\right]\operatorname{dVol}(x)\ .$ (6.1)
Let us prove that the atlas given in the proof of Proposition 3.14 is oriented. A simple computation shows that
@@
Instead of computing the determinant of $A$, we will compute the determinant of the matrix $A^{\prime}$, which coincides with $A$ in all columns except the first one, whose sign is reversed (note that in this case, $\det(A)=-\det(A^{\prime})$). In particular, we have
$A^{\prime}_{ij}=\frac{\|u\|^{2}\delta_{ij}-2u_{i}u_{j}}{\|u\|^{4}}\ .$
The matrix $A^{\prime}$ is diagonalizable. Its eigenvalues are $\frac{1}{\|u\|^{2}}$ with multiplicity $d-1$ and $-\frac{1}{\|u\|^{2}}$ with multiplicity $1$, with eigenvectors
@@
$=\begin{pmatrix}\frac{2(1+\|u\|^{2})\delta_{1,i}-4u_{i}u_{1}}{(1+\|u\|^{2})^{2}}&\cdots&\frac{2(1+\|u\|^{2})\delta_{d,i}-4u_{i}u_{d}}{(1+\|u\|^{2})^{2}}&\frac{4u_{i}}{(1+\|u\|^{2})^{2}}\end{pmatrix}\ ,$
@@
where $\chi(d)$ is the square root of a $\chi^{2}$ random variable with $d$ degrees of freedom. Let us show that $f(s)$ and $D^{i}f(s)$ are uncorrelated, and thus independent (the result is standard for random fields with $\kappa\in C^{2}$, but our assumptions here are more general). From Lemma 5.2 we obtain
@@
$\mathbb{E}\left[\mathcal{H}_{d-1}(A_{u}(T))\right]=$ $\frac{e^{-u^{2}/2}}{\sqrt{2\pi}}\int_{\mathbb{S}^{d}}\mathbb{E}\left[\sqrt{D^{1}T(x)^{2}+\cdots+D^{d}T(x)^{2}}\right]\operatorname{dVol}(x)$
@@ 7   Numerical evidence
Our first experiment is concerned with the fractal class. In this case, the expected value of the boundary length of an excursion set is infinity, because its Hausdorff dimension is strictly larger than $1$. Of course, the numerical experiments will only produce finite values, because the boundary length is computed using a finite pixelization on the sphere. However we manage to illustrate the theoretical behavior by simply increasing the resolution of the experiment: as expected the numerical approximation of the boundary length grows without limit as the pixelization becomes finer and finer. This is in stark contrast with the behavior in the regular class, see Figure 1. The effect of depth is illustrated in Figure 1 where the finite order approximation of the $1$-Hausdorff measure grows larger and larger as the depth increases.
@@
This way, one can cover the low-disorder ($a=1$), sparse ($a=1+\sqrt{2}$) and high-disorder ($a=9$) regimes. Figure 2 is in accordance with Theorem 3.12; the length of the nodal lines exhibits different behaviors depending on the activation function: it diverges exponentially to 00 as the depth increases in the low-disorder case, it remains constant in the sparse case, and it diverges to $\infty$ in the high-disorder scenario. To prevent underflow phenomena in the low-disorder case ($a=1$), we have only considered values of $L$ less than $10$. We note that even though the value of $L$ is very small, we already observe very small values for the length.
In high-disorder cases $(a=9)$, we observe that the simulation and the theoretical value given in Theorem 3.12 are very close for $L<40$. However, for larger $L$ the simulation values are much lower. This behavior is due to the fact that, for very large $L$, the function lives on increasingly higher frequencies (see [24] for more details). For computational reasons, our experiments were conducted using frequencies up to $1536$, and thus, for sufficiently large $L$, we inevitably lose information. This is confirmed by Figure 3, which shows the percentage of variance explained by using the first $1536$ frequencies: as can be seen, the explained variance starts to decrease precisely at $L=40$.
@@ Acknowledgements
Appendix A The relationship between spectral index and covariance regularity
In this appendix, we explore the relationship between the spectral index (cfr. Definition 3.2) and covariance regularity (cfr. Definition 3.1). Since the covariance function can be expressed in terms of Gegenbauer polynomials (2.1), we first recall their definition and prove an integral representation.
@@ A.1   Gegenbauer polynomials
[…]
@@
The first step of our proof is to give an integral representation for the Gegenbauer polynomials. To estabilish this result, we will use the following formula ([47], eq. (4.10.3)), which holds for Jacobi polynomials with $\alpha=\beta=\lambda-1/2$:
@@
The result below has some independent interest and may be known already, but we failed to locate any references and hence we give a full proof.
@@
[…]
@@
[…]
@@
[…]
@@
To prove the next two lemma, we use the polylogarithmic function. For any fixed $s>1$, we define
@@
$\operatorname{Li}_{s}(z)=\Gamma(1-s)(\ln z^{-1})^{s-1}+\sum_{k=0}^{\infty}\frac{\zeta(s-k)}{k!}(\ln z)^{k},\;\;$ $s\not\in\mathbb{N}\text{ and }|\ln z|<2\pi$ (A.2)
@@
$\sum_{\ell=1}^{\infty}\ell^{-s}G_{\ell;d}(\cos(\theta))$ =$=\begin{cases}\zeta(s)-K_{1}\theta^{s-1}+o(\theta^{s-1})&\text{ if }1<s<3\ ,\\ \zeta(s)-K_{2}\theta^{2}+K_{3}\theta^{s-1}+o(\theta^{s-1})&\text{ if }3<s<5\ ,\\ \zeta(s)-K_{2}\theta^{2}+K_{4}\theta^{4}+o(\theta^{4})&\text{ if }s\geq 5\ ,\end{cases}$
@@
using the integral representation of Gegenbauer polynomials (cf.  Lemma A.1) we have
@@
Using Lemma A.5, the previous equation becomes, as $\theta\to 0^{+}$,
@@
  • Let $31<s<$ 3, using the previous identity and since
@@
    indeed, $\sin((s-1)/2\pi)>0$, $\Gamma(1-s)$ and $Q_{0,s-1,d}>0$ (see Lemma A.5 below).
  • Let $53<s<$ 5. A similar computation shows that
@@
$\sum_{\ell=1}^{\infty}\ell^{-s}G_{\ell;d}(\cos(\theta))=\begin{cases}\zeta(2)-K_{5}\theta+K_{6}\theta^{2}\ln(\theta)+o(\theta^{2}\ln(\theta))&\text{ if }s=2\ ,\\ \zeta(3)-K_{7}\theta^{2}|\ln\theta|+o(\theta^{2}\ln\theta)&\text{ if }s=3\ ,\\ \zeta(4)-K_{2}\theta^{2}+K_{8}\theta^{3}+o(\theta^{3})&\text{ if }s=4\ ,\\ \zeta(5)-K_{2}\theta^{2}+K_{9}\theta^{4}\ln(\theta)+o(\theta^{4}\ln\theta)&\text{ if }s=5\ ,\\ \zeta(s)-K_{2}\theta^{2}+K_{4}\theta^{4}+o(\theta^{4})&\text{ if }s\geq 6\ .\end{cases}$
@@
  • If $s-1$ is even, then from (A.4) we obtain
@@
    $k!C_{k}=\begin{cases}{(-1)^{k/2}}\zeta(s-k)&\text{ if }k\text{ even and }k\neq s-1\\ (-1)^{(s-1)/2}H_{s-1}&\text{ if }k=s-1\text{ even}\\ 0&\text{ otherwise }\ ,\end{cases}$
@@
    Using Lemma A.5 and Lemma A.6 we have,
@@
    Using Lemma A.5 and Lemma A.6, we obtain
@@
A.2   Proof of Proposition 3.4
Using the expansion (2.1) of the covariance function in terms of Gegenbauer polynomials, the definition of $n_{\ell,d}$ and the form of the angular power spectrum in Definition 3.2, we have
@@
$1-\kappa(\cos(\theta))$ $=\sum_{\ell=1}^{\infty}\sum_{s=\alpha+1}^{\alpha+d+N}b_{s}\ell^{-s}(1-G_{\ell,d}(\cos(\theta)))\ .$
Exchanging the sums, we obtain
$1-\kappa(\cos(\theta))=\sum_{s=\alpha+1}^{\alpha+d+N}b_{s}Q_{s}(\theta)\ ,$
where
$Q_{s}(\theta)=\sum_{\ell=1}^{\infty}\ell^{-s}(1-G_{\ell,d}(\cos(\theta)))\ .$
@@
$\mathbb{E}[|T(x)-T(y)|^{2}]=D_{1}\theta_{x,y}^{\alpha}+o(\theta_{x,y}^{\alpha})\qquad$ $\alpha\in[0,2)$
$\mathbb{E}[|T(x)-T(y)|^{2}]=D_{2}\theta^{2}_{x,y}+D_{3}\theta_{x,y}^{\alpha}+o(\theta_{x,y}^{\alpha})\qquad$ $\alpha\in(2,4)\ ,$
$\mathbb{E}[|T(x)-T(y)|^{2}]=D_{4}\theta^{2}_{x,y}+D_{5}\theta^{4}_{x,y}\log(\theta_{x,y})+o(\theta_{x,y}^{4}\log(\theta_{x,y}))\qquad$ $\alpha=4\ ,$
$\mathbb{E}[|T(x)-T(y)|^{2}]=D_{4}\theta^{2}_{x,y}+D_{6}\theta^{4}_{x,y}+o(\theta_{x,y}^{4})\qquad$ $\alpha>4,\ .$

### v2

@@ 5.1 Proof of item i)
$\lim_{h,k\to 0}A_{h,k}$ where (5.4)
$A_{h,k}$ $=A_{h,k}(s,s)=\frac{1}{hk}\mathbb{E}\!\left[\big(f(s+he_{i})-f(s)\big)\big(f(s+ke_{i})-f(s)\big)\right].$
@@
$\Sigma_{f_{N}}(s,t)=\mathbb{E}[T(\varphi_{N}^{-1}(s))T(\varphi_{N}^{-1}(t))]=\kappa\left(1-\frac{2\|s-t\|^{2}}{(1+\|s\|^{2})(1+\|t\|^{2})}\right)\ .$ (5.5)
Set
$E(x,y)=\frac{2(x-y)^{2}}{(1+\|s+xe_{i}\|^{2})(1+\|s+ye_{i}\|^{2})}\,.$
Since for every $x,y$, $E(x,y)\geq 0$ and $1-E(x,y)\to 1$ as $x,y\to 0$, there exists $\delta>0$ such that
$\frac{1}{2}\leq 1-E(x,y)\leq 1\qquad\forall(x,y)\in B(0,3\delta)\ ,$
where $B(u,r)$ denotes the open ball in $\mathbb{R}^{2}$ with center $u\in\mathbb{R}^{2}$ and radius $r$. Thus, the function
$G:\,[-\delta,\delta]^{2}\to[-1,1]\qquad G(x,y)=\kappa(1-E(x,y))$
is well defined. Let $h,k\in\mathbb{R}$ such that $|h|,|k|\leq\delta$. Then
$A_{h,k}=\frac{G(h,k)-G(h,0)-G(0,k)+G(0,0)}{hk}\,.$
For every $k$, we define
$\phi_{hk}:\,[0,1]\to\mathbb{R}\qquad t\mapsto G(h,tk)-G(0,tk)\ .$
Then
$A_{h,k}=\frac{\phi_{hk}(1)-\phi_{hk}(0)}{hk}\ .$
Since $\kappa\in C^{1}([-1,1])$ and $E\in C^{\infty}(\mathbb{R}^{2})$, we have $\phi_{hk}\in C^{1}([0,1])$ and hence
$\phi_{hk}(1)-\phi_{hk}(0)=\int_{0}^{1}\phi_{hk}^{\prime}(t)\mathrm{d}t=k\int_{0}^{1}\Big(\partial_{2}G(h,tk)-\partial_{2}G(0,tk)\Big)\mathrm{d}t\ .$
We set
$\psi_{y}(t):[-\delta,\delta]\to\mathbb{R}\qquad t\mapsto\partial_{2}G(t,y)\ .$
In Lemma˜5.3 we prove that $\psi_{y}$ is Lipschitz and therefore absolutely continuous. Hence its derivative exists almost everywhere, and it is therefore almost everywhere equal to $\partial_{12}G(\cdot,y)$. Then
$A_{h,k}=\frac{1}{h}\int_{0}^{1}\int_{0}^{h}\partial_{12}G(r,tk)\mathrm{d}r\mathrm{d}t=\int_{0}^{1}\int_{0}^{1}\partial_{12}G(rh,tk)\mathrm{d}r\mathrm{d}t\ .$
Using (5.11) and (5.12) (with $y=0$), by dominated convergence we obtain
$\lim_{h,k\to 0}A_{h,k}=\frac{4\kappa^{\prime}(1)}{(1+\|s\|^{2})^{2}}\,.$ (5.6)
We have thus proved i).
Lemma 5.3.
The function $\psi_{y}$ is Lipschitz for every $y$.
Proof.
A simple computation shows that
$\partial_{12}G(x,y)=$ $\kappa^{\prime\prime}(1-E(x,y))\partial_{1}E(x,y)\partial_{2}E(x,y)-\kappa^{\prime}(1-E(x,y))\partial_{12}E(x,y)\,.$ (5.7)
Now, $E\in C^{\infty}(\mathbb{R}^{2})$ and $\kappa\in C^{1}([-1,1])$, thus there exists $M_{1}>0$ such that
$\sup_{x,y\in[-\delta,\delta]}\Big|\kappa^{\prime}(1-E(x,y))\partial_{12}E(x,y)\Big|=M_{1}\,.$ (5.8)
Define
$M(x,y)=\kappa^{\prime\prime}(1-E(h+y,y))\partial_{1}E(x,y)\partial_{2}E(x,y)\ .$
A trivial computation shows that, as $h\to 0$,
$M(y+h,y)=-\frac{16h^{2}\kappa^{\prime\prime}(1-E(h+y,y))}{(1+\|s+ye_{i}\|^{2})^{2}(1+\|s+(h+x)e_{i}\|^{2})}+O(h^{3})\,.$
Hence, using the definition of CRI,
$M(y+h,y)=$ $-\frac{16c_{1}\beta(\beta-1)h^{2}}{(1+\|s+(h+y)e_{i}\|^{2})^{2}(1+\|s+ye_{i}\|^{2})^{2}}\times$
$\quad\times\Bigg[\Bigg(\frac{2h^{2}}{(1+\|s+(y+h)e_{i}\|^{2})(1+\|s+ye_{i}\|^{2})}\Bigg)^{\beta-2}\hskip-10.0pt+o(h^{\beta-2})\Bigg]+O(h^{3})$
$=O\Big(h^{2(\beta-1)\wedge 3}\Big)\,.$
Since $\beta>1$, there exists $\varepsilon>0$ such that
$\sup_{x,y\in[-\delta,\delta]\atop{|x-y|<\varepsilon}}|M(x,y)|<1\,.$ (5.9)
Otherwise, for $|x-y|>\varepsilon$, $E(x,y)<1$. Since $\kappa\in C^{2}((-1,1))$, there exists $M_{2}>0$ such that
$\sup_{x,y\in[-\delta,\delta]\atop{|x-y|\geq\varepsilon}}|M(x,y)|=M_{2}\,.$ (5.10)
Combining (5.7) with (5.8), (5.9) and (5.10), we get
$\sup_{x,y\in[-\delta,\delta]}\Big|\partial_{12}G(x,y)\Big|\leq M_{1}+(M_{2}+1)=M\,.$ (5.11)
The previous computation shows also that $\psi^{\prime}_{y}(x)$ is well-defined and coincides with $\partial_{12}G(x,y)$ for every $x\neq y$. Moreover,
$\lim_{x\to y}\psi_{y}^{\prime}(x)=-\kappa^{\prime}(1)\partial_{12}E(x,y)\Big|_{x=y}=\frac{4\kappa^{\prime}(1)}{(1+\|s+ye_{i}\|^{2})^{2}}\leq M_{1}\,.$ (5.12)
This implies that $\psi_{y}$ is $M$-Lipschitz. ∎
5.2 Covariance computation
Let $s,t\in\mathbb{R}^{d}$ and $i,j\in 1,\dotsc,d$. For every $h,k\in\mathbb{R}$, define
$A_{h,k}(s,t;i,j)=\mathbb{E}\left[\frac{f(t+he_{i})-f(t)}{h}\cdot\frac{f(s+ke_{j})-f(s)}{k}\right]\ .$
Note that this definition coincides with the one given in (5.4), in the case $s=t$ and $i=j$.
@@ 5.3 Proof of item ii)
The second term in (5.16) can be written, using Lemma˜5.2 and (5.15), as
@@
Finally to compute the third term in (5.16) one can use Lemma˜5.2, eq.˜5.14 and the formula for $\kappa^{\prime}$ obtained by deriving (3.1) to get that
@@ 5.4 Proof of item iii)
as a consequence, taking $k=d$ we have $2d\rho(\beta)\leq d+d$ and therefore (5.3) holds with $\zeta_{1}=2d$, $\zeta_{2}=d$ and $K=C^{k}(2k-1)!!$.
From (5.3), we can compute $\mathbb{E}[D^{i}f(s)^{2}]$ and $\mathbb{E}[D^{i}f(t)^{2}]$, whence to estimate (5.17) we only need to evaluate $\mathbb{E}[D^{i}f(s)D^{i}f(t)]$. Using again Lemma˜5.2 we obtain
@@
Since $1<\beta<2$ we have also $\|t-s\|^{2}=o(\|t-s\|^{2\beta-2})$ and therefore
@@
6  Proof of Theorem˜3.14: boundary volumes in the Kac-Rice class
In Proposition˜6.2, we compute the expected values of the boundary volumes of a $C^{1}$ isotropic Gaussian random field. Theorem˜3.14 follows using the fact that $\kappa_{L}^{\prime}(1)=\kappa^{\prime}(1)^{L}$ as proved in  [27]. We note that, using Proposition˜3.16, under the assumptions of Theorem˜3.14, the fields $T_{L}$ are $C^{1}$ a.s for every $L$.
In the proof of Proposition˜6.2, we use the following result that can be obtained combining Theorem .2 and Theorem 2.2 in [7].
@@
$\mathbb{E}[\mathcal{H}_{d-1}(X^{-1}(u))]=\int_{\mathbb{S}^{d}}\mathbb{E}\left.\big[\|\nabla X(t)\|\,\right|\,X(t)=u\big]p_{X(t)}(u)\operatorname{dVol}(t)\ ,$
@@
$\mathbb{E}\left[\mathcal{H}_{d-1}(T^{-1}(u))\right]=\omega_{d-1}\kappa^{\prime}(1)^{1/2}e^{-u^{2}/2}\qquad a.s.\ .$
@@
Theorem˜6.1 implies
$\mathbb{E}\left[\mathcal{H}_{d-1}(T^{-1}(u))\right]=\int_{\mathbb{S}^{d}}\mathbb{E}\left[\sqrt{\nabla T(x)^{t}\nabla T(x)}\;|\;T(x)=u\right]p_{T(x)}(u)\operatorname{dVol}(x)\ .$
@@
$\mathbb{E}\left[\mathcal{H}_{d-1}(T^{-1}(u))\right]=\frac{1}{\sqrt{2\pi}}e^{-{u^{2}}/2}\int_{\mathbb{S}^{d}}\mathbb{E}\left[\sqrt{\nabla T(x)^{t}\nabla T(x)}\;|\;T(x)=u\right]\operatorname{dVol}(x)\ .$ (6.1)
Let us prove that the atlas given in the proof of Proposition˜3.16 is oriented. A simple computation shows that
@@
[…]
$\widetilde{A}_{ij}=\frac{\|u\|^{2}\delta_{ij}-2u_{i}u_{j}}{\|u\|^{4}}\ .$
The matrix $\widetilde{A}$ is diagonalizable. Its eigenvalues are $\frac{1}{\|u\|^{2}}$ with multiplicity d1d-1 and $\frac{1}{\|u\|^{2}}$ with multiplicity $1$, with eigenvectors
@@
$=\begin{pmatrix}\frac{2(1+\|u\|^{2})\delta_{1,i}-4u_{i}u_{1}}{(1+\|u\|^{2})^{2}}&\cdots&\frac{2(1+\|u\|^{2})\delta_{d,i}-4u_{i}u_{d}}{(1+\|u\|^{2})^{2}}&\frac{4u_{i}}{(1+\|u\|^{2})^{2}}\end{pmatrix}\$,
@@
where $\chi(d)$ is the square root of a $\chi^{2}$ random variable with $d$ degrees of freedom. Let us show that $f(s)$ and $D^{i}f(s)$ are uncorrelated, and thus independent (the result is standard for random fields with $\kappa\in C^{2}$, but our assumptions here are more general). From Lemma˜5.2 we obtain
@@
$\mathbb{E}\left[\mathcal{H}_{d-1}(T^{-1}(u))\right]=$ $\frac{e^{-u^{2}/2}}{\sqrt{2\pi}}\int_{\mathbb{S}^{d}}\mathbb{E}\left[\sqrt{D^{1}T(x)^{2}+\cdots+D^{d}T(x)^{2}}\right]\operatorname{dVol}(x)$
@@ 7  Numerical evidence
Our first experiment is concerned with the fractal class. In this case, the expected value of the boundary length of an excursion set is infinity, because its Hausdorff dimension is strictly larger than $1$. Of course, the numerical experiments will only produce finite values, because the boundary length is computed using a finite pixelization on the sphere. However we manage to illustrate the theoretical behavior by simply increasing the resolution of the experiment: as expected the numerical approximation of the boundary length grows without limit as the pixelization becomes finer and finer. This is in stark contrast with the behavior in the regular class, see Figure˜1. The effect of depth is illustrated in Figure˜1 where the finite order approximation of the $1$-Hausdorff measure grows larger and larger as the depth increases.
@@
This way, one can cover the low-disorder ($a=1$), sparse ($a=1+\sqrt{2}$) and high-disorder ($a=9$) regimes. Figure˜2 is in accordance with Theorem 3.14; the length of the nodal lines exhibits different behaviors depending on the activation function: it diverges exponentially to 0 as the depth increases in the low-disorder case, it remains constant in the sparse case, and it diverges to $\infty$ in the high-disorder scenario. To prevent underflow phenomena in the low-disorder case ($a=1$), we have only considered values of $L$ less than $10$. We note that even though the value of $L$ is very small, we already observe very small values for the length.
In high-disorder cases $(a=9)$, we observe that the simulation and the theoretical value given in Theorem˜3.14 are very close for $L<40$. However, for larger $L$ the simulation values are much lower. This behavior is due to the fact that, for very large $L$, the function lives on increasingly higher frequencies (see [27] for more details). For computational reasons, our experiments were conducted using frequencies up to $1536$, and thus, for sufficiently large $L$, we inevitably lose information. This is confirmed by Figure˜3, which shows the percentage of variance explained by using the first $1536$ frequencies: as can be seen, the explained variance starts to decrease precisely at $L=40$.
@@ Acknowledgements
Appendix A The relationship between spectral index and CRI
In this appendix, we explore the relationship between the spectral index (cfr. Definition˜3.2) and covariance regularity (cfr. Definition˜3.1). Since the covariance function can be expressed in terms of Gegenbauer polynomials (2.1), we first recall their definition and prove an integral representation.
@@ A.1 Gegenbauer polynomials
    $\int_{-1}^{1}P_{\ell}^{(\alpha,\beta)}(x)P_{\ell^{\prime}}^{(\alpha,\beta)}(x)(1-x)^{\alpha}(1+x)^{\beta}\,dx=0,$
@@
The first step of our proof is to give an integral representation for the Gegenbauer polynomials. To establish this result, we will use the following formula ([52], eq. (4.10.3)), which holds for Jacobi polynomials with $\alpha=\beta=\lambda-1/2$:
@@
The result below has some independent interest and may be known already, but we failed to locate any references, and hence we give a full proof.
@@
$G_{\ell,d}(\cos\theta)=C_{d}\frac{-i}{|\sin^{d-2}(\theta)|}\int_{e^{-i\theta}}^{e^{i\theta}}h^{\ell}\left(h^{2}-2h\cos\theta+1\right)^{(d-3)/2}\,dh.$
@@
$G_{\ell,d}(\cos\theta)=\frac{C_{d}}{|\sin^{d-2}(\theta)|}\int_{-\theta}^{\theta}e^{i\ell\psi}\left(e^{i2\psi}-2e^{i\psi}\cos\theta+1\right)^{(d-3)/2}e^{i\psi}\,d\psi.$
@@
$G_{\ell,d}(\cos\theta)=\frac{C_{d}}{\sin^{d-2}(\theta)}\int_{-\theta}^{\theta}e^{i(\ell+(d-1)/2)\psi}\left(2\cos\psi-2\cos\theta\right)^{(d-3)/2}\,d\psi.$
@@
To prove the next two lemmas, we use the polylogarithmic function. For any fixed $s>1$, we define
@@
$\operatorname{Li}_{s}(z)=\Gamma(1-s)(\ln z^{-1})^{s-1}+\sum_{k=0}^{\infty}\frac{\zeta(s-k)}{k!}(\ln z)^{k},\;\;\$$s\not\in\mathbb{N}\text{ and }|\ln z|<2\pi$ (A.2)
@@
$\sum_{\ell=1}^{\infty}\ell^{-s}G_{\ell;d}(\cos(\theta))$ $=\begin{cases}\zeta(s)-K_{1}\theta^{s-1}+o(\theta^{s-1})&\text{ if }1<s<3\ ,\\ \zeta(s)-K_{2}\theta^{2}+K_{3}\theta^{s-1}+o(\theta^{s-1})&\text{ if }3<s<5\ ,\\ \zeta(s)-K_{2}\theta^{2}+K_{4}\theta^{4}+o(\theta^{4})&\text{ if }s\geq 5\ ,\end{cases}$
@@
using the integral representation of Gegenbauer polynomials (see  Lemma˜A.1) we have
@@
Using Lemma˜A.5, the previous equation becomes, as $\theta\to 0^{+}$,
@@
  • Let $1<s<3$, using the previous identity and since
@@
    indeed, $\sin((s-1)/2\pi)>0$, $\Gamma(1-s)$ and $Q_{0,s-1,d}>0$ (see Lemma˜A.5 below).
  • Let $3<s<5$. A similar computation shows that
@@
[…]
@@
  • If s1s-1 is even, then from (A.4) we obtain
@@
[…]
@@
    Using Lemma˜A.5 and Lemma˜A.6 we have,
@@
    Using Lemma˜A.5 and Lemma˜A.6, we obtain
@@
A.2 Proof of Proposition˜3.4
Using the expansion (2.1) of the covariance function in terms of Gegenbauer polynomials, the definition of $n_{\ell,d}$ and the form of the angular power spectrum in Definition˜3.2, we have
@@
$1-\kappa(\cos(\theta))$ $=\sum_{\ell=1}^{\infty}\sum_{s=\alpha+1}^{\alpha+d+N}b_{s}\ell^{-s}(1-G_{\ell,d}(\cos(\theta)))\ .$ (A.5)
We note that, for $s>1$, we have
$\sum_{\ell=1}^{\infty}\ell^{-s}=\zeta(s)<\infty\ .$
Then, using Lemma˜A.3,
$Q_{s}(\theta)=\sum_{\ell=1}^{\infty}\ell^{-s}(1-G_{\ell,d}(\cos(\theta))<\infty\ .$
Since the second sum in (A.5) has a finite number of terms,
$1-\kappa(\cos(\theta))=\sum_{s=\alpha+1}^{\alpha+d+N}b_{s}\sum_{\ell=1}^{\infty}\ell^{-s}(1-G_{\ell,d}(\cos(\theta)))=\sum_{s=\alpha+1}^{\alpha+d+N}b_{s}Q_{s}(\theta)\ .$
@@
$\mathbb{E}[|T(x)-T(y)|^{2}]=D_{1}\theta_{x,y}^{\alpha}+o(\theta_{x,y}^{\alpha})\qquad$ $\alpha\in[0,2)\ ,$
$\mathbb{E}[|T(x)-T(y)|^{2}]=D_{2}\theta_{x,y}^{2}|\ln(\theta_{x,y})|+o(\theta_{x,y}^{2}|\ln(\theta_{x,y})|)\qquad$ $\alpha=2\ ,$
$\mathbb{E}[|T(x)-T(y)|^{2}]=D_{3}\theta^{2}_{x,y}+D_{4}\theta_{x,y}^{\alpha}+o(\theta_{x,y}^{\alpha})\qquad$ $\alpha\in(2,4)\ ,$
$\mathbb{E}[|T(x)-T(y)|^{2}]=D_{5}\theta^{2}_{x,y}+D_{6}\theta^{4}_{x,y}\log(\theta_{x,y})+o(\theta_{x,y}^{4}\log(\theta_{x,y}))\qquad$ $\alpha=4\ ,$
$\mathbb{E}[|T(x)-T(y)|^{2}]=D_{5}\theta^{2}_{x,y}+D_{7}\theta^{4}_{x,y}+o(\theta_{x,y}^{4})\qquad$ $\alpha>4,\ .$
@@
$\kappa(1-t)=\kappa(1)+B_{2}\arccos(1-t)^{2}\ln(\arccos(1-t))$
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}+}+o(\arccos(1-t)^{2}\ln(\arccos(1-t)))\qquad$ $\alpha=2\ ,$